If a car starts from rest and accelerates at 5.0 m/sec2 for 6.0 seconds, how far did it travel and how fast was it moving at the end of 6.0 seconds?
1
Expert's answer
2012-08-30T09:08:46-0400
The distance traveled is

L = at²/2 = 5*6²/2 = 90 m

and the velocity reached is

V = at = 5*6 = 30 m/s.